# Real-World Evaluation of two Cooperative Intersection Management Approaches

### v5

@@ VII Conclusion
In this paper, an optimization-based and an RL-based approach for cooperative maneuver planning in mixed traffic at urban intersections have been evaluated in a realistic simulation and in real-world experiments. Especially in simulated fully automated traffic, cooperative maneuvers achieved a significant improvement of traffic efficiency. Even in simulated mixed traffic with only 50 % CAV penetration, the planning approaches successfully performed cooperative maneuvers in about one third of the scenarios and increased the efficiency, despite not relying on any cooperation from HDVs. Overall, the optimization-based planner yielded higher efficiency gains, while also causing a larger increase of criticality measures. Experiments with three prototype CAVs in public traffic demonstrated the real-world applicability of both cooperative planners and showed a significant efficiency increase. Future work will include the evaluation of further cooperative use cases in urban traffic involving human-driven connected vehicles and vulnerable road users.

### v6

@@ VI-B Evaluation in Real-World
The three real-world CAVs all use different motion planning algorithms that were adapted to demonstrate the cooperative use case and were not yet fully tuned to exactly meet the time constraints resulting from the maneuver planning. Therefore, a criticality metric evaluation on the real-world test drives would not be of high significance, which should, of course, be addressed by further research.
@@ VII Conclusion
In this paper, an optimization-based and an RL-based approach for cooperative maneuver planning in mixed traffic at urban intersections have been evaluated in a realistic simulation and in real-world experiments. Especially in simulated fully automated traffic, cooperative maneuvers achieved a significant improvement of traffic efficiency. Even in simulated mixed traffic with only 50 % CAV penetration, the planning approaches successfully performed cooperative maneuvers in about one third of the scenarios and increased the efficiency, despite not relying on any cooperation from HDVs. Overall, the optimization-based planner yielded higher efficiency gains, while also causing a larger increase of criticality measures. Experiments with three prototype CAVs in public traffic demonstrated the real-world applicability of both cooperative planners and showed a significant efficiency increase. Future work will include the evaluation of further cooperative use cases in urban traffic involving human-driven connected vehicles and vulnerable road users. Moreover, incorporating state uncertainty information from the environment model in the cooperative planning algorithms might be a promising avenue for future research.